# OPENXRD: A Comprehensive Benchmark and Enhancement Framework for LLM/MLLM XRD Question Answering

### v1

@@
[…]
tcb@breakable
[…]
OPENXRD: A Comprehensive Benchmark and Enhancement Framework for LLM/MLLM XRD Question Answering
@@
This work presents OPENXRD, an open-book pipeline designed for crystallography question answering, which integrates textual prompts with concise supporting content generated by GPT-4.5. Instead of using scanned textbooks, which may lead to copyright issues, OPENXRD generates compact, domain-specific references that help smaller models understand key concepts in X-ray diffraction (XRD). We evaluate OPENXRD on a well-defined set of 217 expert-level XRD questions by comparing different vision-language models, including GPT-4 and LLaVA-based frameworks such as Mistral, LLaMA, and QWEN, under both closed-book (without supporting material) and open-book (with supporting material) conditions. Our experimental results show significant accuracy improvements in models that use the GPT-4.5-generated summaries, particularly those with limited prior training in crystallography. OPENXRD uses knowledge from larger models to fill knowledge gaps in crystallography and shows that AI-generated texts can help smaller models reason more effectively in scientific tasks. While the current version of OPENXRD focuses on text-based inputs, we also explore future extensions such as adding real crystal diagrams or diffraction patterns to improve interpretation in specialized materials-science contexts. Overall, OPENXRD shows that specialized open-book systems can be useful in materials science and provides a foundation for broader natural language processing (NLP) tools in critical scientific fields.
footnotetext:  These authors contributed equally to this work.footnotetext: a Department of Computer Science, University of Rochester, Rochester, New York 14627, USA.footnotetext: b Department of Mechanical Engineering, University of Rochester, Rochester, 14627, NY, USA. footnotetext: c Material Science Program, University of Rochester, Rochester, 14627, NY, USA. footnotetext: d Staff Scientist, Laboratory for Laser Energetics (LLE), University of Rochester, Rochester, 14627, NY, USA. footnotetext: † The codes and benchmarking dataset can be accessed at https://github.com/niaz60/OpenXRD. Project webpage: https://niaz60.github.io/OpenXRD/.
@@ 1 Introduction
Crystallography is the scientific discipline concerned with determining the arrangement of atoms and molecules in crystalline solids 1, 2, 3. This structural understanding is crucial for elucidating material properties, such as symmetry, geometry, and physical characteristics 4, 5. This understanding is crucial for progressing materials science, especially in areas like metallurgy, pharmaceuticals, and semiconductor technology 6, 7, 8, 9, 10. Central to crystallography is X-ray diffraction (XRD), a key experimental technique that enables researchers to uncover detailed information about crystal structures by examining how X-rays interact with crystalline lattices 11, 12, 13.
@@
The foundational principle of XRD is Bragg’s Law, discovered by Lawrence Bragg in 1912 14, 15. Bragg’s law describes how the angle of a diffracted X-ray beam depends on the wavelength of the X-rays and the spacing between atoms and molecules within the material 16, 17. Using XRD, researchers analyze crystal structures to determine important parameters such as atomic arrangements, phase composition, unit cell size, grain size, crystallinity, strain, and lattice defects 18, 19, 20, 21, 22, 23.
@@
Advances in the analysis of large XRD data have demonstrated that conventional deep learning approaches, such as Convolutional Neural Networks (CNNs) and Graph Neural Networks (GNNs), are powerful tools, particularly for identifying or predicting crystal parameters space group labels 24, 25, 26, lattice constants 27, 28, or phase compositions 29, 30. These models achieve remarkable numerical accuracy, demonstrating their efficacy for crystal classification and structural analysis tasks.
@@
However, despite their success in quantitative predictions, these deep-learning techniques are limited in providing interpretative and explanatory insights into the underlying physics or chemistry of XRD data 31, 32, 33. This lack of interpretability restricts their ability to offer meaningful explanations about the material properties or structural characteristics encoded within the diffraction data. This lack of interpretability restricts the broader application of deep-learning approaches in crystallographic research, where understanding underlying phenomena is essential 34, 35, 36.
@@
In contrast, recent progress in Natural Language Processing (NLP) driven by Large Language Models (LLMs), such as GPT-based architectures, has transformed various computational domains 37, 38, 39. LLMs demonstrate exceptional capabilities in open-ended question answering 40, 41, multi-step reasoning 42, 43, domain-specific question answering 44, 45, 46, predicting material properties 47, 48, 49, and extracting information from complex datasets 50, 51, 52. These powerful language models can potentially bridge the interpretability gap faced by traditional deep-learning models when applied to scientific fields, including materials science and crystallography 53, 54, 55.
@@
For instance, Antunes et al. introduced CrystaLLM, an autoregressive LLM trained on millions of Crystallographic Information Files (CIFs), which encode detailed crystallographic data, including atomic coordinates, symmetry operations, lattice parameters, and space group information, to generate plausible crystal structures from given chemical compositions, demonstrating the potential of text-based crystal structure generation 56. In addition, Johansen et al. introduced deCIFer, an autoregressive language model on powder XRD data to produce complete crystal structures in CIF format, attaining high match rates between generated structures and experimental diffraction profiles 57.
@@
Further extending these capabilities, Choudhary introduced DiffractGPT, a generative pretrained transformer that predicts atomic structures directly from XRD patterns, particularly when chemical information is provided 58. AtomGPT, another transformer-based model by Choudhary, effectively predicts material properties such as formation energies, electronic bandgaps, and superconducting transition temperatures with accuracy comparable to GNNs 59. Beyond generative tasks, LLMs have also been employed for scientific question-answering and knowledge retrieval in the materials domain. For example, the LLaMP framework combines an LLM with a materials database to fetch and reason over crystallographic data, reducing hallucinations and improving factual accuracy in materials science QA 51.
@@
In this study, we investigate whether providing domain-specific context generated by a stronger model, such as supporting textual material, can significantly enhance the performance of weaker models or alternative configurations of the same model on specialized XRD-related questions. To rigorously test this hypothesis, we constructed a carefully selected dataset consisting of 217 multiple-choice XRD questions, each reviewed and approved by a domain expert at the Ph.D. level. Each question has only one correct answer among four provided choices, covering a range of topics from fundamental principles to complex scenarios, including issues such as basic structural geometry, unit cell dimensions, and coordination environments, to more complex concepts, including fundamental equations and symmetry analysis. Initially, we evaluate the performance of various Large LLMs, including GPT-4.5, GPT-4, O1, O3, and so forth, under closed-book conditions, where models rely solely on their pretrained internal knowledge without external assistance. Our initial results show that GPT-4.5 clearly outperforms other tested models. Based on this outcome, we employ GPT-4.5 to generate approximately one-page textual summaries. These summaries intentionally avoid explicitly stating correct answers yet offer adequate context and guidance to facilitate correct reasoning. We then reassess the performance of the other models under open-book conditions to quantify the improvement when provided with these contextual summaries, aiming to quantify the improvement enabled by this supplementary textual material.
@@ 2 Methods
2.1 Dataset selection and model processing
To evaluate language model performance in crystallography, we curated a domain-specific multiple-choice question-answering dataset, comprising 217 items focused exclusively on X-ray diffraction and crystallography. To construct this dataset, we partially drew on content from Cullity’s Introduction to X-ray Diffraction 11. Each question in the dataset includes a concise prompt, a set of three to four answer options, a single correct answer, a brief explanation providing the rationale, and a corresponding subtask label. These subtasks span a range of crystallographic domains, including crystal structure fundamentals, diffraction geometry, unit cell analysis, and related topics. As illustrated in Figure 1, the distribution of subtask labels highlights the diversity and complexity of the topics covered in our dataset, ranging from basic definitions to advanced structural phenomena. All questions were reviewed and validated by a domain expert to ensure scientific correctness and comprehensive coverage of essential XRD subtasks.
Fig.  1: A word cloud of the subtask labels from our human-curated crystallography dataset. Larger words indicate subtasks with a higher number of questions, illustrating the breadth of topics (e.g., diffraction fundamentals, lattice geometry, advanced structure analysis). This distribution underscores the diversity and domain complexity captured in our QA benchmark, spanning both basic definitions (e.g., counting crystal systems) and intricate reflections (e.g., twin boundaries, space-group anomalies).
@@ 2.1 Dataset selection and model processing
Fig.  2: The expert-reviewed version provides essential clarifications about the angular dependence of atomic scattering factors, emphasizes their direct relationship to atomic electron count, and highlights practical implications for interpreting diffraction data, elements often missed in AI-generated explanations.
Question What happens if the path difference between scattered rays is $n\lambda$? Options
(A) They cancel each other
(B) They reinforce each other …
(C) etc.
Expert-Reviewed Supporting Material “When two waves differ in path by $n\lambda$ (where $n$ is an integer), they are perfectly in phase and interfere constructively, resulting in amplitude reinforcement. This is a fundamental principle in wave interference and forms the basis of the Bragg diffraction condition in crystallography.”
Fig.  3: An open-book mode example with expert-reviewed supporting material that clearly explains how a path difference of $n\lambda$ leads to constructive interference, with additional contextual information about its relevance to crystallography. The correct answer is highlighted in green.
@@
The OPENXRD evaluation framework is designed to assess the reasoning capabilities of language models in crystallography under two distinct conditions: closed-book and open-book modes.
In the closed-book mode, the model receives only the question along with its corresponding multiple-choice options. It must rely entirely on its pre-trained internal knowledge to answer, without access to any additional resources or clarifications. This setting evaluates the model’s intrinsic understanding of crystallographic principles and its ability to perform domain-specific reasoning based solely on its learned representations.
In contrast, the open-book mode supplements the question and answer choices with a brief, domain-relevant textual passage. This supporting material contains essential background information or clarifications pertinent to the question, similar to consulting a short textbook excerpt. The aim is to evaluate the model’s ability to integrate external context with its internal knowledge in order to reason more effectively.
Together, these two evaluation settings provide complementary insights. While the closed-book mode tests baseline domain knowledge and internal consistency, the open-book mode examines the model’s capacity to synthesize and apply external scientific information. This dual-mode structure is particularly valuable for identifying the limits of a model’s independent expertise versus its ability to benefit from curated guidance.
@@
(Open-Book, GPT-4.5 Supporting Knowledge)
@@
Fig.  4: Illustration of our open-book QA pipeline for crystallography. In closed-book mode, the model sees only the question (left rotated box). In open-book mode, it also receives domain-specific supporting textual material (right rotated box), which is concatenated and fed to the LLM (center pipeline), producing the final QA result.
@@ 2.3 Supporting Textual Material Generation
2.3.1 AI-Generated Supporting Materials (Open-Book, Without Expert Review)
@@
The generation process followed three key principles. First, each passage summarized fundamental crystallographic concepts relevant to the specific question. Second, care was taken to avoid directly stating the correct answer, ensuring that models must still reason rather than extract solutions verbatim. Third, the length and content were controlled to maintain clarity and focus, typically resulting in brief paragraphs spanning half to one page.
@@ 2.3.2 AI-Generated Materials with Expert Review
We engaged three PhD students specializing in crystallography to review and refine the supporting materials. These experts, with four to seven years of research experience in XRD, crystal structure analysis, and materials characterization, were instructed to correct any technical inaccuracies in AI-generated materials, improve explanations with precise domain terminology, ensure comprehensive coverage of relevant concepts, improve clarity while maintaining concision, remove potentially misleading information, and add critical contextual details missing from the original materials. As demonstrated in Figure 2, the expert-reviewed version provides substantial improvements over the initial AI-generated explanations. Furthermore, Figure 3 illustrates a representative example from the multiple-choice question-answering dataset, showcasing the original question, corresponding answer choices, and the AI-generated explanation as revised by an expert.

### v2

# OPENXRD: A Comprehensive Benchmark Framework for LLM/MLLM XRD Question Answering

@@
OPENXRD: A Comprehensive Benchmark Framework for
LLM/MLLM XRD Question Answering
@@
We introduce OPENXRD, a comprehensive benchmarking framework for evaluating large language models (LLMs) and multimodal LLMs (MLLMs) in crystallography question answering. The framework measures context assimilation, or how models use fixed, domain-specific supporting information during inference. The framework includes 217 expert-curated X-ray diffraction (XRD) questions covering fundamental to advanced crystallographic concepts, each evaluated under closed-book (without context) and open-book (with context) conditions, where the latter includes concise reference passages generated by GPT-4.5 and refined by crystallography experts. We benchmark 74 state-of-the-art LLMs and MLLMs, including GPT-4, GPT-5, O-series, LLaVA, LLaMA, QWEN, Mistral, and Gemini families, to quantify how different architectures and scales assimilate external knowledge. Results show that mid-sized models (7B–70B parameters) gain the most from contextual materials, while very large models often show saturation or interference and the largest relative gains appear in small and mid-sized models. Expert-reviewed materials provide significantly higher improvements than AI-generated ones even when token counts are matched, confirming that content quality, not quantity, drives performance. OPENXRD offers a reproducible diagnostic benchmark for assessing reasoning, knowledge integration, and guidance sensitivity in scientific domains, and provides a foundation for future multimodal and retrieval-augmented crystallography systems.
footnotetext: These authors contributed equally to this work.footnotetext: a Dept. of Computer Science, University of Rochester, Rochester, New York 14627, USA.footnotetext: b Dept. of Mechanical Engineering, University of Rochester, Rochester, 14627, NY, USA. footnotetext: c Material Science Program, University of Rochester, Rochester, 14627, NY, USA. footnotetext: d Staff Scientist, Laboratory for Laser Energetics (LLE), University of Rochester, Rochester, 14627, NY, USA. footnotetext: † The codes and benchmarking dataset can be accessed at
https://github.com/niaz60/OpenXRD.
Project webpage: https://niaz60.github.io/OpenXRD/.
@@ 1 Introduction
Crystallography is the scientific discipline concerned with determining the arrangement of atoms and molecules in crystalline solids 37, 87, 97. This structural understanding is crucial for elucidating material properties, such as symmetry, geometry, and physical characteristics 100, 64. This understanding is crucial for progressing materials science, especially in areas like metallurgy, pharmaceuticals, and semiconductor technology 86, 4, 13, 1, 41. Central to crystallography is X-ray diffraction (XRD), a key experimental technique that enables researchers to uncover detailed information about crystal structures by examining how X-rays interact with crystalline lattices 25, 3, 15.
@@
The foundational principle of XRD is Bragg’s Law, discovered by Lawrence Bragg in 1912 16, 14. Bragg’s law describes how the angle of a diffracted X-ray beam depends on the wavelength of the X-rays and the spacing between atoms and molecules within the material 82, 91. Using XRD, researchers analyze crystal structures to determine important parameters such as atomic arrangements, phase composition, unit cell size, grain size, crystallinity, strain, and lattice defects 30, 47, 99, 38, 83, 28.
@@
Advances in the analysis of large XRD data have demonstrated that conventional deep learning approaches, such as Convolutional Neural Networks (CNNs) and Graph Neural Networks (GNNs), are powerful tools, particularly for identifying or predicting crystal parameters space group labels 89, 109, 78, lattice constants 33, 18, or phase compositions 48, 93. These models achieve remarkable numerical accuracy, demonstrating their efficacy for crystal classification and structural analysis tasks.
@@
However, despite their success in quantitative predictions, these deep-learning techniques are limited in providing interpretative and explanatory insights into the underlying physics or chemistry of XRD data 29, 92, 46. This lack of interpretability restricts their ability to offer meaningful explanations about the material properties or structural characteristics encoded within the diffraction data. This lack of interpretability restricts the broader application of deep-learning approaches in crystallographic research, where understanding underlying phenomena is essential 103, 66, 19.
@@
In contrast, recent progress in Natural Language Processing (NLP) driven by Large Language Models (LLMs), such as GPT-based architectures, has transformed various computational domains 53, 36, 26. LLMs demonstrate exceptional capabilities in open-ended question answering 108, 45, multi-step reasoning 20, 79, domain-specific question answering 107, 12, 59, predicting material properties 24, 88, 84, and extracting information from complex datasets 50, 21, 35. These powerful language models can potentially bridge the interpretability gap faced by traditional deep-learning models when applied to scientific fields, including materials science and crystallography 106, 90, 101.
@@
For instance, Antunes et al. introduced CrystaLLM, an autoregressive LLM trained on millions of Crystallographic Information Files (CIFs), which encode detailed crystallographic data, including atomic coordinates, symmetry operations, lattice parameters, and space group information, to generate plausible crystal structures from given chemical compositions, demonstrating the potential of text-based crystal structure generation 10. In addition, Johansen et al. introduced deCIFer, an autoregressive language model on powder XRD data to produce complete crystal structures in CIF format, attaining high match rates between generated structures and experimental diffraction profiles 44.
@@
Further extending these capabilities, Choudhary introduced DiffractGPT, a generative pretrained transformer that predicts atomic structures directly from XRD patterns, particularly when chemical information is provided 23. AtomGPT, another transformer-based model by Choudhary, effectively predicts material properties such as formation energies, electronic bandgaps, and superconducting transition temperatures with accuracy comparable to GNNs 22. Beyond generative tasks, LLMs have also been employed for scientific question-answering and knowledge retrieval in the materials domain. For example, the LLaMP framework combines an LLM with a materials database to fetch and reason over crystallographic data, reducing hallucinations and improving factual accuracy in materials science QA 21.
@@
In this study, we investigate whether providing domain-specific context generated by a stronger model, such as supporting textual material, can significantly enhance the performance of weaker models or alternative configurations of the same model on specialized XRD-related questions. To rigorously test this hypothesis, we constructed a carefully selected dataset consisting of 217 multiple-choice XRD questions, each reviewed and approved by a domain expert at the Ph.D. level. Each question has only one correct answer among four provided choices, covering a range of topics from fundamental principles to complex scenarios, including issues such as basic structural geometry, unit cell dimensions, and coordination environments, to more complex concepts, including fundamental equations and symmetry analysis. Initially, we evaluate the performance of various Large LLMs, including GPT-4.5, GPT-4, O1, O3, and so forth, under closed-book conditions, where models rely solely on their pretrained internal knowledge without external assistance. Our initial results show that GPT-4.5 clearly outperforms other tested models. Based on this outcome, we employ GPT-4.5 to generate approximately one-page textual summaries. These summaries intentionally avoid explicitly stating correct answers yet offer adequate context and guidance to facilitate correct reasoning. We then reassess the performance of the other models under open‑book conditions to quantify the improvement when provided with these contextual summaries, aiming to quantify the improvement enabled by this supplementary textual material.
It is important to clarify that OPENXRD is a context-assimilation benchmark rather than a retrieval system, and as such is complementary to Retrieval-Augmented Generation (RAG) rather than competing with it. While RAG is a deployment architecture that couples a retriever with a generator to fetch answer-bearing passages from large corpora at inference time, OPENXRD deliberately removes the retrieval confound by providing fixed, curated, answer-guiding passages that avoid revealing correct answers. Accordingly, our open-book setting should be interpreted as an oracle (“Gold-Standard”) RAG condition: the generator receives perfectly relevant, expert-verified context (i.e., perfect retrieval), so the measured accuracy is an upper bound on what an end-to-end RAG system could achieve with the same base model under ideal retrieval. This design isolates the language model’s ability to integrate external guidance from confounding factors such as retrieval quality, chunking, and ranking. OPENXRD does not perform retrieval and does not aspire to be a deployable QA system; instead, it provides a controlled, reproducible setting to disclose how language models react to guidance—when it helps, when it distracts, and how sensitivity varies with model family, token budget, and content quality. These are properties that end-to-end RAG often obscure because retrieved passages can directly contain the answer. The same evaluation harness can benchmark RAG systems by replacing our oracle helper passages with retrieved chunks from a crystallography corpus, thereby decomposing end-to-end RAG accuracy into retrieval quality and language model assimilation capability. Thus, OPENXRD serves as a diagnostic complement for RAG research, not a competing alternative.
@@ 2 Methods
2.1 Dataset Selection and Curation
To evaluate language model performance in crystallography, we curated a domain-specific multiple-choice question-answering dataset comprising 217 expert-reviewed items focused exclusively on X-ray diffraction and crystallographic reasoning. To construct this dataset, we partially drew on content from Cullity’s Introduction to X-ray Diffraction 25. Each question includes a concise prompt, three to four answer options, a single correct answer, a brief explanation providing the scientific rationale, and a associated subtask label that enables fine-grained performance analysis across different crystallographic domains. The dataset encompasses 81 distinct crystallographic subtasks, covering a broad span of domain knowledge including crystal structure fundamentals, diffraction geometry, reciprocal-space concepts, Bragg-law applications, peak indexing, space-group interpretation, unit-cell analysis, scattering physics, microstructure effects, and additional specialized topics (see Figure 1 and Figure 2). We adopt a multiple-choice format in this foundational benchmark because it enables unambiguous, reproducible scoring across many models, avoiding the evaluator subjectivity that arises in open-ended generation. This design choice is intended to isolate context assimilation effects under controlled conditions rather than to fully represent real-world crystallographic workflows. Evaluating open-ended, problem-solving tasks (e.g., short-answer explanations and multi-step derivations) is an important direction but is beyond the scope of the present benchmark and is discussed as future work.
Fig. 1: Example question from the OPENXRD dataset in closed-book format. Each question includes a concise prompt, multiple-choice options (3-4 answers), with one correct answer. This illustrates the baseline evaluation condition where models must rely solely on their internal knowledge without external supporting materials.
As illustrated in Figure 2, the distribution of subtask labels highlights the diversity and complexity of the topics covered in our dataset, ranging from basic definitions to advanced structural phenomena. All questions were reviewed and validated by a domain expert to ensure scientific correctness and comprehensive coverage of essential XRD subtasks.
Fig. 2: A word cloud of the subtask labels from our human-curated crystallography dataset. Larger words indicate subtasks with a higher number of questions, illustrating the breadth of topics (e.g., diffraction fundamentals, lattice geometry, advanced structure analysis). This distribution underscores the diversity and domain complexity captured in our QA benchmark, spanning both basic definitions (e.g., counting crystal systems) and intricate reflections (e.g., twin boundaries, space-group anomalies).
@@ 2.1 Dataset Selection and Curation
Fig. 3: The expert-reviewed version provides essential clarifications about the angular dependence of atomic scattering factors, emphasizes their direct relationship to atomic electron count, and highlights practical implications for interpreting diffraction data, elements often missed in AI-generated explanations.
Fig. 4: An open-book mode example with expert-reviewed supporting material that clearly explains how a path difference of $n\lambda$ leads to constructive interference, with additional contextual information about its relevance to crystallography. The correct answer is highlighted in green.
@@
The OPENXRD evaluation framework is designed to assess the reasoning capabilities of language models in crystallography under two distinct conditions: closed-book and open-book modes. In the closed-book mode, the model receives only the question along with its corresponding multiple-choice options. It must rely entirely on its pre-trained internal knowledge to answer, without access to any additional resources or clarifications. This setting evaluates the model’s intrinsic understanding of crystallographic principles and its ability to perform domain-specific reasoning based solely on its learned representations. In contrast, the open-book mode supplements the question and answer choices with a brief, domain-relevant textual passage. This supporting material contains essential background information or clarifications pertinent to the question, similar to consulting a short textbook excerpt. Because this passage is fixed and expert-reviewed (not retrieved), open-book results estimate the generator-side ceiling under perfect retrieval (an oracle / Gold-Standard RAG scenario) rather than the performance of a full retriever+generator pipeline. The aim is to evaluate the model’s ability to integrate external context with its internal knowledge in order to reason more effectively. Together, these two evaluation settings provide complementary insights. While the closed-book mode tests baseline domain knowledge and internal consistency, the open-book mode examines the model’s capacity to synthesize and apply external scientific information. This dual-mode structure is particularly valuable for identifying the limits of a model’s independent expertise versus its ability to benefit from curated guidance.
@@
(Open‑Book, GPT-4.5 Supporting Knowledge)
@@
Fig. 5: Illustration of our open-book QA pipeline for crystallography. In closed-book mode, the model sees only the question (left rotated box). In open-book mode, it also receives domain-specific supporting textual material (right rotated box), which is concatenated and fed to the LLM (center pipeline), producing the final QA result.
@@ 2.3 Supporting Textual Material Generation
2.3.1 AI-Generated Supporting Materials (Open‑Book, Without Expert Review)
@@
The generation process followed three key principles. First, each passage summarized fundamental crystallographic concepts relevant to the specific question. Second, care was taken to avoid directly stating the correct answer, ensuring that models must still reason rather than extract solutions verbatim. Third, the length and content were controlled to maintain clarity and focus, typically resulting in brief paragraphs spanning half to one page. Our standardized generation protocol was: Given each question and answer choices (without revealing which option is correct), we prompted GPT-4.5 (temperature=0.7, top_p=0.95, max_tokens=800) with: “Given the following question: [QUESTION] and answer choices: [OPTIONS], provide a concise explanation of [RELEVANT CONCEPT] that would help answer this question without revealing the correct answer. Focus on fundamental principles, key definitions, and relevant physical mechanisms. Target length: 500–600 tokens.” Passages that revealed answers or contained technical errors were regenerated with revised prompts emphasizing guidance without revelation.
@@ 2.3.2 AI-Generated Materials with Expert Review
We engaged three PhD students specializing in crystallography to review and refine the supporting materials. These experts, with four to seven years of research experience in XRD, crystal structure analysis, and materials characterization, were instructed to correct any technical inaccuracies in AI-generated materials, improve explanations with precise domain terminology, ensure comprehensive coverage of relevant concepts, improve clarity while maintaining concision, remove potentially misleading information, and add critical contextual details missing from the original materials.
As demonstrated in Figure 3, the expert-reviewed version provides substantial improvements over the initial AI-generated explanations. Furthermore, Figure 4 illustrates a representative example from the multiple-choice question-answering dataset, showcasing the original question, corresponding answer choices, and the AI-generated explanation as revised by an expert.
2.3.3 Quality Dimensions for Expert Review
To make expert refinement reproducible, reviewers applied four explicit quality dimensions to all 217 passages: (1) Technical accuracy—terminology, equations, and physical principles must be correct and precise; (2) Conceptual precision—the targeted mechanism must be distinguished from related but distinct phenomena to avoid distraction-based errors; (3) Pedagogical structure—content is organized from definitions to mechanisms to question-specific implications; (4) Contextual relevance—remove tangential information and focus on the knowledge gap the question targets. Reviewers kept token counts within 5% of the AI-generated versions to preserve token-matched comparisons.